Calculate and Use Overhead Rate

Selected data for the consulting department of Austin Consulting, Inc., follow:

Estimated consulting overhead cost for the year $810,000
Estimated direct labor cost for the year (@ $9/hr.) 450,000
Actual manufacturing overhead cost for January 48,000
Actual direct labor cost for January (1,200 hours) 33,000

Assuming that direct labor cost is the basis for applying consulting overhead,


a. Calculate the predetermined overhead rate.

_________%

b. Prepare a journal entry that applies consulting overhead for January.

General Journal
Description Debit Credit
Answer Answer Answer
Answer Answer Answer

c. By what amount is consulting overhead over- or underapplied in January?

Consulting overhead was ________ by $_______ in January.

Preparing a Job Order Cost Sheet

Riverwood Accounting Company has the following account in its cost records:

Work in Process—Jones Audit
Direct labor 120,000 Services completed 253,518
Project overhead 162,000

Riverwood applies overhead to projects at a predetermined rate based on direct labor costs. Assume that Riverwood uses a job order costing system and that Jones Audit is the only job in process at the end of the period. Complete the following cost sheet for services still in process for Jones Audit.

Job Order Cost Sheet
Jones Audit (Services in Process)
Direct labor Answer
Project overhead Answer
Total cost Answer
Answer #1

1a) Predetermine overhead rate = 810000/450000 = 180% of labor cost

b. Prepare a journal entry that applies consulting overhead for January.

General Journal
Description Debit Credit
Work in process (33000*1.80) 59400
Consulting overhead 59400

c. By what amount is consulting overhead over- or underapplied in January?

Consulting overhead was Overapplied by (59400-48000) = 11400 in January.

2) Calculate following

Complete the following cost sheet for services still in process for Jones Audit.

Job Order Cost Sheet
Jones Audit (Services in Process)
Direct labor (28482/2.35) 12120
Project overhead (12120*1.35) 16362
Total cost 28482